OpenStudy (anonymous):

Verify 2sin(x)=cot(x)[cos(x)tan(x)+sin(x)]

OpenStudy (anonymous):

cot(x) times tan(x) = 1 so, the first part would be just cosine x. cot(x) * sin(x) = cos(x)/sin(x) * sin(x)/1 = cos(x) So, you have 2sin(x)=cos(x)+cos(x) which simplifies to 2sin(x)=2cos(x) sin(x)=cos(x) thus it's not an identity.

OpenStudy (anonymous):

They are only equal at 45 degrees.